OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4 a. 2 b. 3 c. All real numbers d. No solution

OpenStudy (anonymous):

-2x-6= -2x-2 -4 -2x-6=-2x-6 0=0

OpenStudy (anonymous):

c= all real numbers